1. A die is rolled 300 times; it lands five 58 times. Is this evidence significant enough to conclude that the die is not fai I need answer part 2
Answer #1

ANSWER:

Ans:

1)sample proportion

=58/300

=0.1933

Test statistic:

z=(0.1933-0.1667)\sqrt{}(0.1667*(1-0.1667)/300)

z=1.239

critical z value=+/-1.96

Fail to reject the null hypothesis.

There is not sufficient evidnece to conclude that the die is not fairly balanced.

2)

lower cut off=0.1667-1.96*\sqrt{(0.1667*(1-0.1667)/300)

=0.1245

upper cut off=0.1667-1.96*\sqrt{(0.1667*(1-0.1667)/300)

=0.2089

when true p=0.25

z(0.1245)=(0.1245-0.25)\sqrt{(0.25*(1-0.25)/300)

=-5.02

z(0.2089)=(0.2089-0.25)\sqrt{(0.25*(1-0.25)/300)

=-1.645

P(type II error)=P(-5.02<z<-1.646)

=P(z<-1.646)-P(z<-5.02)

=0.0498-0.0000

=0.0498